when light goes from water (n=1.333) into air (n=1.000), the angle of a fraction is 25.1. What is the angle of incidence


1
Expert's answer
2021-05-13T09:21:16-0400

We can find the angle of incidence from the Snell's Law:


"n_isin\theta_i=n_rsin\theta_r,""\theta_i=sin^{-1}(\dfrac{n_rsin\theta_r}{n_i}),""\theta_i=sin^{-1}(\dfrac{1.0\cdot sin25.1^{\circ}}{1.333})=17.6^{\circ}."